OpenStudy (anonymous):

A baseball is thrown up into the air. The ball's height in feet is modeled by the equation h(t) = -16t2 + 32t + 10. (h = height, t = time in seconds) Find the maximum height of the ball.

OpenStudy (anonymous):

magic formula for vertex is \[-\frac{b}{2a}\]

OpenStudy (anonymous):

in this case \[a=-16,b=32,-\frac{b}{2a}=1\]

OpenStudy (anonymous):

replace h by 1 to get your answer. i get \[-16+32+10=16+10=26\]